OpenStudy (anonymous):

You have made 3 out of 10 of your last free throws. How many free throws can you expect to make if you shoot 100 free throws? a) 30 b) 50 c) 60 d) 75

OpenStudy (anonymous):

a) 30, do you need to know how?

OpenStudy (anonymous):

Sure Why Not Thx :D

OpenStudy (anonymous):

Since 10 free throws were attempted, and 100 are now being attempted, 100 divided by 10 is 10, so multiply the 3 by 10 to get 30 free throws out of 100

OpenStudy (anonymous):

It makes sense once you get used to it

OpenStudy (anonymous):

Yes It sure does ! :D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

I get it

OpenStudy (anonymous):

There is another thing called a proportion, it's a bit more complicated, but also is easy once you get used to it. 3 free throws made out of 10 can be a fraction, which is \[\frac{ 3 }{ 10 }\]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

That 1 way cool

OpenStudy (anonymous):

Since at first you do not know how many free throws were made out of 100, let x represent the value you are trying to find, which is \[\frac{ x }{ 100 }\]

OpenStudy (anonymous):

Both fractions are equal

OpenStudy (anonymous):

Cool

OpenStudy (anonymous):

Thank You So Much SO Kind, Go Explanation Good Job Thx Love It :D * ~ * ^ _ ^!!!

OpenStudy (anonymous):

So \[\frac{ 3 }{ 10 }\] = \[\frac{ x }{ 100 }\]

OpenStudy (anonymous):

There is one more step to that

OpenStudy (anonymous):

Whats That????????

OpenStudy (anonymous):

There is a thing called "cross multiplying" and that would be multiplying 10 and x, and 3 and 100. 10 multiplied by x is 10x, and 3 multiplied by 100 is 300. This would be \[\frac{ 10x }{ 10 }\] = \[\frac{ 300 }{ 10 }\] 300/10 means divide, so divide both to get 30

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

If your teacher hasn't taught you it yet you can surprise her

OpenStudy (anonymous):

LOL, Your right

OpenStudy (anonymous):

You think you could help me with maybe with 3 more?

OpenStudy (anonymous):

Suppose you write the letters M-I-S-S-I-S-S-I-P-P-I on cards and mix them in a hat. You choose one card from the hat without looking. What is the probability of picking a P? a) 2/11 b) 4/11 c) 4/7 d) 7/4

OpenStudy (anonymous):

With 1 sorry

OpenStudy (anonymous):

if not thats find

OpenStudy (anonymous):

It is 2/11

OpenStudy (anonymous):

11 letters, 2 are P